# Predicting New Research Directions in Materials Science using Large Language Models and Concept Graphs

### v1

@@ Discussion
The code used to produce the results of this paper is publicly available at: https://github.com/aimat-lab/materials_concepts. It is licensed under the GPL 3.0+ License.
@@ Acknowledgements
Creating a chemical element parser from scratch is hard. We thank Chris Konop for an outline of how a chemical formula can be recursively defined. This made our implementation way more straightforward. We acknowledge support by the Federal Ministry of Education and Research (BMBF) under Grant No. 01DM21001B (German-Canadian Materials Acceleration Center). We acknowledge funding by the German Research Foundation (DFG) under Germany’s Excellence Strategy via the Excellence Cluster “3D Matter Made to Order” (3DMM2O, EXC-2082/1–390761711). Funding by the German Research Foundation (DFG) through the collaborative research center CRC 1249 N-Heteropolycycles as Functional Materials (SFB 1249, Project C13) and through DFG Projekt 436506789 is gratefully acknowledged. This work was partly carried out with the support of the Karlsruhe Nano Micro Facility (KNMFi, www.knmf.kit.edu), a Helmholtz Research Infrastructure at Karlsruhe Institute of Technology (KIT, www.kit.edu). We acknowledge funding from the Helmholtz Metadata Collaboration (HMC) through the project AIMWORKS. This work was performed on the HoreKa supercomputer funded by the Ministry of Science, Research and the Arts Baden-Württemberg and by the Federal Ministry of Education and Research. The authors acknowledge support by the state of Baden-Württemberg through bwHPC.
Supplemental Information
Dataset
Filtering
Cleaning
Figure S1: Overview of the dataset generation: (a) fetch publications from OpenAlex venues related to materials sciences, (b) filtering based on title and abstract (c) removal of selected characters.
Chemical Formulas
  1. 1.
    The subsequent word is another element
  2. 2.
    The subsequent word is an integer together with a + or -
  3. 3.
    The subsequent word is an integer, a + or a -
  4. 4.
    The subsequent word is a decimal
  5. 5.
    The subsequent word is a composition of elements of the periodic table
LLM Fine-tuning
Hugging Face’s Python library was used to fine-tune the LLaMa-2-13B model [wolf_huggingfaces_2020, touvron_llama_2023-1]. We prepare our annotated data by constructing a single string per data point as follows <s> + abstract + "#KEYWORDS#" + concepts + </s> where concepts corresponds to the string version of a Python list, e.g. [’brittle epoxy’, ’butadiene nitrile liquid rubber’, ...]. <s> and </s> represent the beginning of sequence (BOS) and, respectively, end of sequence (EOS) token for the LLaMa models. #KEYWORDS# serves as an own custom delimiter to signal the model when concept extraction should start. For the LLaMa tokenizer, we set pad_token_id = 0, bos_token_id = 1, and eos_token_id = 2. Being a CausalLM, we set the labels equal to the input token ids. Our LoraConfig is r = 16 (Lora attention dimension), lora_alpha = 32, target_modules=["q_proj", "v_proj"] (modules where the adapter is applied), and lora_dropout = 0.01 (dropout probability) [dettmers2023qlora]. Training took around an hour on a single A100 GPU.
Hyperparameters of the link prediction models
Table S1: Optimized hyperparameters of link prediction models.[a]
Model Layers Neurons per layer Learning rate
Baseline 7 [20, 300, 180, 108, 64, 10, 1] $5\cdot 10^{-4}$
Concept embeddings 5 [1536, 1024, 819, 10, 1] $1\cdot 10^{-3}$
Combination of features 7 [1556, 1556, 933, 559, 335, 10, 1] $1\cdot 10^{-3}$
Concept extraction
@@ Abstract
Removal of Metal Impurities from the Spent Hydroprocessing Catalyst and Acid Leaching Kinetics of Aluminum. The impurity metal of Al in spent hydroprocessing catalyst MoNi/Al2O3 was removed by HCl, and the metals of Mo and Ni were preliminarily enriched. The results show that the leaching efficiency of Al was 88.62%, and the leaching efficiency of Mo and Ni were 16.32% and 28.74%, respectively. The results were achieved under optimal leaching conditions: the particle size was 150 m, the concentration of HCl was 4 mol/L, the leaching temperature was, and the leaching time was 120 min. The kinetics of the leaching behavior of Al showed that the acid leaching reaction of Al was in accorded with the equation 1-2/3/X-1-X 2/3/ =K 2 t, R 2 =0.97734, which was controlled by internal diffusion. X-ray diffraction analysis of the leaching residue revealed the existence of residual metals MoMoO3 , MoO2ClO42 and NiNiS. The separation of Al from Mo and Ni has been preliminarily realized, which is conducive to further efficient recovery of Mo and Ni.
LLM Concepts Rake Concepts
acid leaching kinetic aluminium acid leaching aluminium impurity aluminium removal hydroprocessing catalyst hydroprocessing catalyst impurity metal impurity metal impurity removal spent hydroprocessing catalyst x ray diffraction analysis acid leaching kinetics acid leaching reaction al showed efficient recovery equation xx impurity metal internal diffusion xray diffraction analysis leaching behavior leaching efficiency leaching efficiency leaching residue revealed leaching temperature leaching time metal impurities optimal leaching conditions particle size preliminarily enriched preliminarily realized residual metals momoo mooclo results show spent hydroprocessing catalyst spent hydroprocessing catalyst monialo
Table S2: Concept Extraction for W3027426609
@@
LLM Concepts Rake Concepts
anchor effect carbon fiber reinforcement composite strengthening mechanism elongation grain refinement in situ interfacial reaction interface property interfacial reaction load transfer mechanical property nano sized phase short carbon fiber strengthening mechanism tensile property ultimate tensile strength yield strength anchor effect c sf c sf content c sf content ultimate tensile strength c sf increases c sf reinforced metal matrix composites composites increases different contents elongation decreases slightly elongation reaches grain refinement interfacial bonding state interfacial reaction investigated according investigated interface load transfer load transfer main strengthening mechanism matrix forms mechanical properties mechanical properties strengthening mechanisms nb phase embedded intimately nb phases nbnbsi composites reinforced short carbon fiber c sf short carbon fiber reinforced nbnbsi composites situ interfacial reaction strengthening behaviors strengthening mechanisms studied interfacial nanosized nb phase forms tensile properties valuable information yield strength
Table S3: Concept Extraction for W3192967258. This demonstrates the ability of LLMs to extract concepts in a normalizing manner.
@@
LLM Concepts Rake Concepts
catia flow of plastic gas assisted injection molding gas assisted molded parts gas channel design gas channel dimension gas fingering gas injected plate moldflow numerical simulation physical simulation design simulated parts different dimensions gas assisted injection moldinggaim gas channels gas fingering gasassisted injected plate gasassisted molded parts gaschannel design mm mm gas channels numerical simulating experiment catia numerical simulation optimized proportion range physical simulating experiment physical simulation plate gas channels plate thickness results show semicircular gas channel semicircular gas channels square gas channels structure design
Table S4: Concept Extraction for W2375662644
@@
LLM Concepts Rake Concepts
agglomerated binding force al2o3 coating coating microstructure dry sliding condition friction and wear tester functional structure lamellar structure partially melted region phase composition plasma spraying scanning electron microscopy wear property wear resistance wear tester x ray diffraction dry sliding condition fullymelted region fullymelted regions show function structure alo tio functional structure lamellar structure low agglomerated binding force mainly ascribe micro powder coating ml friction nano powder coating nano powder coating consists nano powder coating resulted nano powders partiallymelted region partiallymelted regions retain phase compositions plasma sprayed powders structure scanning electron microscopy steel functional structure using micro wear properties wear resistance wear test wear tester xray diffraction respectively results show
Table S5: Concept Extraction for W2521598419
@@
LLM Concepts Rake Concepts
amorphous alloy avrami analysis crystallization behavior crystallization process devitrification behavior differential scanning calorimetry dsc eutectic crystallization fully amorphous ribbon inhomogeneity isothermal dsc experiment kinetic analysis kissinger analysis melt spinning metastable phase non isothermal dsc experiment partially amorphous ribbon polymorphic transformation primary crystallization quenching device ribbon length ribbon thickness structural state transmission electron microscopy x ray diffraction al followed alsm amorphous alloys various alxsmx alloys alsm display eutectic crystallization amorphous phases avrami analyses careful control crystallization behaviour crystallization processes different mechanisms distinct parts dsc traces involve several overlapping peaks kinetic analysis kinetic parameters derived meltspun ribbons metastable intermetallic alsm metastable mixture metastable phases alsm transforms polymorphically nonisothermal dsc experiments partially amorphous ribbons particularly sensitive quenching conditions except quenching device rapidly solidified ribbon thickness singleroller melt spinning structural state subsequent devitrification behaviour thinnest ribbons transformations occurring xray diffraction transmission electron microscopy
Table S6: Concept Extraction for W2047249259
@@
Successive ionic layer adsorption and reaction SILAR trend for nanocrystalline mercury sulfide thin films growth. Mercury sulfide HgS nanocrystalline thin films have been grown onto amorphous glass substrate by successive ionic layer adsorption and reaction SILAR trend at room temperature 27C. The optimized preparative parameters including ion concentration, number of immersion cycles, and pH of the solution are used for fine nanocrystalline film growth. A further study has been made for the structural, surface morphological, optical and electrical properties of the films by using X-ray diffraction XRD, scanning electron microscopy SEM, transmission electron microscopy TEM, optical absorption and dc two point probe method. The as-deposited grown HgS nanocrystalline films exhibited cubic phase, with optical band gap Eg of 2.0eV and electrical resistivity of the order of 103cm. SEM and TEM images confirmed films of smooth surface morphology and nanocrystaline in nature with fine crystallites of 2030nm diameter, respectively.
LLM Concepts Rake Concepts
amorphous glass substrate cubic phase dc two point probe method electrical property electrical resistivity mercury sulfide nanocrystaline mercury sulfide thin films nanocrystaline mercury sulfide thin films growth optical band gap optical property scanning electron microscopy sem successive ionic layer adsorption and reaction successive ionic layer adsorption and reaction trend transmission electron microscopy x ray diffraction dc two point probe method diameter respectively electrical properties electrical resistivity fine crystallites fine nanocrystalline film growth grown onto amorphous glass substrate immersion cycles optical band gap eg reaction silar trend room temperature c smooth surface morphology structural surface morphological optical successive ionic layer adsorption tem images confirmed films
Table S7: Concept Extraction for W1993498556: ’and’ conjunction is not misinterpreted. Extraction leaves out the abbreviation.
Concept graph
Table S8: 25 most frequent concepts with their occurrences and degree in the concept graph.
@@
Figure S2: Distribution of node degrees. The node degree of 35 is the most common, occurring 2637 times. The concept with the highest degree is ’mechanical property’. The concept with the lowest degree is ’flexible pneumatic cylinder’. Both axes are in log scale.
Figure S3: Concept Centralization over different years. The graph illustrates the cumulative percentage of edges connected to nodes when sorted by their degree. A steep curve indicates that a small percentage of nodes (highly central nodes) are connected to a large percentage of edges. The years 1980, 1990, 2000, 2010, and 2020 are compared to show how concept centralization increases over time in our Concept Graph.
Table S9: 5 examplary concepts and their respective 5 nearest neighbors
@@
Link prediction performance
Table S10: Confusion matrix classes for the link predictions of the baseline and Concept Embeddings model split by previous node distances
Positives Negatives
Model $d_{\text{prev}}$ TP FN FP TN
Baseline 2 212 78 115,807 0,751,021
3 001 16 005,342 1,124,056
4 000 00 000,000 0,003,467
Concept embeddings 2 203 87 120,575 0,746,253
3 006 11 026,209 1,103,189
4 000 00 000,027 0,003,440
Report generation
Figure S4: Schematic representation of the report structure. ’Own’ ($C_{\text{own}}$) and ’Other’ ($C_{\text{other}}$) concepts are respectively represented in yellow and blue.
@@
[…]
||| $\deg(c_{{\text{other}}})<=100\And 3<=\text{dist}(c_{\text{own}},c_{{\text{other}}}<=6)\}\quad.$
To avoid unspecific suggestions, i.e. broad terms like ‘materials science’, we exclude concepts based on their degree in the concept graph $G$. To prevent obvious, e.g. synonyms, and too far-fetched concepts, we integrate semantic knowledge by using the euclidean distance in the map of the materials science 5 that is a two dimensional projection of the obtained concept embeddings in Concept extraction. Note that the values 3 and 6 were arbitrarily chosen after examining the map of materials science.
@@
$S_{\text{(many own)}\times\text{other}}$ :=$:=\{(c_{{\text{other}}},C_{\text{own},\text{matching}})\in C_{{\text{other}}}\times\{C_{\text{own}}\times...\times C_{\text{own}}\}|C_{\text{own},\text{matching}}$
:=$:=\{c\in C_{\text{own}}|\text{score}(c_{{\text{other}}},c)>p\}\}$
@@
We are interested in the capabilities of LLMs to assist humans by pre-selecting from a broader list of concept combinations. For this, we query two LLMs, GPT-4 and Claude 3 Opus (access May/June 2024), to select three items from a list of combinations and to write an additional paragraph on why this combination is promising. Three combinations are selected from $S_{\text{own}\times\text{own}}\cup S_{\text{own}\times\text{other}}$ and three additional combinations are selected from $S^{\text{filtered}}_{\text{own}\times\text{other}}$, making up a total of six suggested combinations per LLM per report. Note that this report section does not contain new suggestions but rather selects and elaborates on combinations already present in the report. Apart from highlighting particular suggestions, the LLM approach helps to bridge the gap between two concepts by providing additional information on how they could be combined. This is especially helpful for $S_{\text{own}\times\text{other}}$ or $S^{\text{filtered}}_{\text{own}\times\text{other}}$ where the second concept might be unfamiliar to a researcher.
You are an advanced AI agent who has expertise in the realm of the materials sciences. Your goal is to assist researchers looking for promising new concept combinations that will - once investigated - advance mankind.
@@
Figure S5: Prompt for the selection of 3 concept combinations from report sections 1 and 2
@@
Figure S6: Prompt for the selection of 3 concepts combinations from report section 4
@@
Figure S7: Example of an LLM output containing three topic combinations and the corresponding abstracts
Rated Concept Combinations
Table S11: Concept combinations ranked during the interviews (25 examples per category)
@@
B ferroelectric domain interface bonding 3

### v2

@@ Methods
The code used to produce the results of this study is publicly available and permanently archived on Zenodo [47]. The source code is licensed under the GPL 3.0+ License and the active development repository can be found at https://github.com/aimat-lab/materials_concepts.
@@ Acknowledgements
Creating a chemical element parser from scratch is hard. We thank Chris Konop for an outline of how a chemical formula can be recursively defined. This made our implementation way more straightforward. We acknowledge support by the Federal Ministry of Education and Research (BMBF) under Grant No. 01DM21001B (German-Canadian Materials Acceleration Center) (P.F.). We acknowledge funding by the German Research Foundation (DFG) under Germany’s Excellence Strategy via the Excellence Cluster “3D Matter Made to Order” (3DMM2O, EXC-2082/1–390761711) (P.F.). Funding by the German Research Foundation (DFG) through the collaborative research center CRC 1249 N-Heteropolycycles as Functional Materials (SFB 1249, Project C13) and through DFG Projekt 436506789 is gratefully acknowledged (P.F.). This work was partly carried out with the support of the Karlsruhe Nano Micro Facility (KNMFi, www.knmf.kit.edu), a Helmholtz Research Infrastructure at Karlsruhe Institute of Technology (KIT, www.kit.edu) (B.B.). We acknowledge funding from the Helmholtz Metadata Collaboration (HMC) through the project AIMWORKS (P.F.). This work was performed on the HoreKa supercomputer funded by the Ministry of Science, Research and the Arts Baden-Württemberg and by the Federal Ministry of Education and Research. The authors acknowledge support by the state of Baden-Württemberg through bwHPC (P.F.).
Supplementary Information
S1 Concept extraction
Evaluating LLM-based Concept Extraction
Supplementary Fig. 1: Precision, Recall and F1-Score of our LLM-based approach, keyBERT, RAKE and SciSpacy; evaluated against a held-out test set of 5 annotated abstracts, including 138 ground-truth concepts in total.
To construct a held-out test set, three annotators (n=3) manually annotated five abstracts that were not part of the training data. The resulting inter-annotator agreement, measured as the mean pairwise agreement, was 0.345 (pairwise scores: 0.215, 0.326, 0.493; all values rounded to three decimal places).
Importantly, none of the human-annotated concepts were incorrect. Disagreement primarily stemmed from varying emphasis on annotation breadth versus depth. As all extracted concepts were valid, we defined the union of all annotations as the ground truth for evaluating our LLM-based concept extraction approach against established methods, namely KeyBERT, RAKE, and SciSpacy.
RAKE achieved slightly higher recall than the LLM-based approach, primarily due to its extensive extraction of candidate terms. However, this also resulted in substantially lower precision, as for every relevant concept, approximately five irrelevant ones were produced (precision $\approx 16.7\%$). When balancing precision and recall, the LLM-based method outperformed all other approaches, achieving a substantially higher precision at similar recall compared to RAKE.
Notably, concepts produced by the LLM that were not counted as true positives tended to be semantically relevant but either overly broad or overly specific (e.g., ‘aerospace part thickness measurement’), whereas competing methods frequently generated entirely irrelevant outputs such as ‘6148 10 3 cycles’, ‘increase’, ‘interdisciplinary field involving principles’, or ‘317’. This highlights the practical advantage of the LLM-based approach: it substantially reduces the need for manual post-processing, as the extracted concepts are largely free of irrelevant entries.
To enable researchers to benchmark their own models, all labeled data can be found on our GitHub repository.
Examples of extracted concepts
@@ Abstract
Removal of Metal Impurities
from the Spent Hydroprocessing Catalyst and Acid Leaching Kinetics of Aluminum. The impurity metal of Al in spent hydroprocessing catalyst MoNi/Al2O3 was removed by HCl, and the metals of Mo and Ni were preliminarily enriched. The results show that the leaching efficiency of Al was 88.62%, and the leaching efficiency of Mo and Ni were 16.32% and 28.74%, respectively. The results were achieved under optimal leaching conditions: the particle size was 150 m, the concentration of HCl was 4 mol/L, the leaching temperature was, and the leaching time was 120 min. The kinetics of the leaching behavior of Al showed that the acid leaching reaction of Al was in accorded with the equation 1-2/3/X-1-X 2/3/ =K 2 t, R 2 =0.97734, which was controlled by internal diffusion. X-ray diffraction analysis of the leaching residue revealed the existence of residual metals MoMoO3 , MoO2ClO42 and NiNiS. The separation of Al from Mo and Ni has been preliminarily realized, which is conducive to further efficient recovery of Mo and Ni.
LLM Concepts Rake Concepts acid leaching kinetic aluminium acid leaching aluminium impurity aluminium removal hydroprocessing catalyst hydroprocessing catalyst impurity metal impurity metal impurity removal spent hydroprocessing catalyst x ray diffraction analysis acid leaching kinetics acid leaching reaction al showed efficient recovery equation xx impurity metal internal diffusion xray diffraction analysis leaching behavior leaching efficiency leaching efficiency leaching residue revealed leaching temperature leaching time metal impurities optimal leaching conditions particle size preliminarily enriched preliminarily realized residual metals momoo mooclo results show spent hydroprocessing catalyst spent hydroprocessing catalyst monialo
Supplementary Fig. 2: Concept Extraction for W3027426609
@@
LLM Concepts Rake Concepts anchor effect carbon fiber reinforcement composite strengthening mechanism elongation grain refinement in situ interfacial reaction interface property interfacial reaction load transfer mechanical property nano sized phase short carbon fiber strengthening mechanism tensile property ultimate tensile strength yield strength anchor effect c sf c sf content c sf content ultimate tensile strength c sf increases c sf reinforced metal matrix composites composites increases different contents elongation decreases slightly elongation reaches grain refinement interfacial bonding state interfacial reaction investigated according investigated interface load transfer load transfer main strengthening mechanism matrix forms mechanical properties mechanical properties strengthening mechanisms nb phase embedded intimately nb phases nbnbsi composites reinforced short carbon fiber c sf short carbon fiber reinforced nbnbsi composites situ interfacial reaction strengthening behaviors strengthening mechanisms studied interfacial nanosized nb phase forms tensile properties valuable information yield strength
Supplementary Fig. 3: Concept Extraction for W3192967258 (normalization of concepts through LLM extraction).
@@
LLM Concepts Rake Concepts catia flow of plastic gas assisted injection molding gas assisted molded parts gas channel design gas channel dimension gas fingering gas injected plate moldflow numerical simulation physical simulation design simulated parts different dimensions gas assisted injection moldinggaim gas channels gas fingering gasassisted injected plate gasassisted molded parts gaschannel design mm mm gas channels numerical simulating experiment catia numerical simulation optimized proportion range physical simulating experiment physical simulation plate gas channels plate thickness results show semicircular gas channel semicircular gas channels square gas channels structure design
Supplementary Fig. 4: Concept Extraction for W2375662644.
@@
LLM Concepts Rake Concepts agglomerated binding force al2o3 coating coating microstructure dry sliding condition friction and wear tester functional structure lamellar structure partially melted region phase composition plasma spraying scanning electron microscopy wear property wear resistance wear tester x ray diffraction dry sliding condition fullymelted region fullymelted regions show function structure alo tio functional structure lamellar structure low agglomerated binding force mainly ascribe micro powder coating ml friction nano powder coating nano powder coating consists nano powder coating resulted nano powders partiallymelted region partiallymelted regions retain phase compositions plasma sprayed powders structure scanning electron microscopy steel functional structure using micro wear properties wear resistance wear test wear tester xray diffraction respectively results show
Supplementary Fig. 5: Concept Extraction for W2521598419.
@@
LLM Concepts Rake Concepts amorphous alloy avrami analysis crystallization behavior crystallization process devitrification behavior differential scanning calorimetry dsc eutectic crystallization fully amorphous ribbon inhomogeneity isothermal dsc experiment kinetic analysis kissinger analysis melt spinning metastable phase non isothermal dsc experiment partially amorphous ribbon polymorphic transformation primary crystallization quenching device ribbon length ribbon thickness structural state transmission electron microscopy x ray diffraction al followed alsm amorphous alloys various alxsmx alloys alsm display eutectic crystallization amorphous phases avrami analyses careful control crystallization behaviour crystallization processes different mechanisms distinct parts dsc traces involve several overlapping peaks kinetic analysis kinetic parameters derived meltspun ribbons metastable intermetallic alsm metastable mixture metastable phases alsm transforms polymorphically nonisothermal dsc experiments partially amorphous ribbons particularly sensitive quenching conditions except quenching device rapidly solidified ribbon thickness singleroller melt spinning structural state subsequent devitrification behaviour thinnest ribbons transformations occurring xray diffraction transmission electron microscopy
Supplementary Fig. 6: Concept Extraction for W2047249259.
@@
Successive ionic layer adsorption and reaction
SILAR trend for nanocrystalline mercury sulfide thin films growth. Mercury sulfide HgS nanocrystalline thin films have been grown onto amorphous glass substrate by successive ionic layer adsorption and reaction SILAR trend at room temperature 27C. The optimized preparative parameters including ion concentration, number of immersion cycles, and pH of the solution are used for fine nanocrystalline film growth. A further study has been made for the structural, surface morphological, optical and electrical properties of the films by using X-ray diffraction XRD, scanning electron microscopy SEM, transmission electron microscopy TEM, optical absorption and dc two point probe method. The as-deposited grown HgS nanocrystalline films exhibited cubic phase, with optical band gap Eg of 2.0eV and electrical resistivity of the order of 103cm. SEM and TEM images confirmed films of smooth surface morphology and nanocrystaline in nature with fine crystallites of 2030nm diameter, respectively.
LLM Concepts Rake Concepts amorphous glass substrate cubic phase dc two point probe method electrical property electrical resistivity mercury sulfide nanocrystaline mercury sulfide thin films nanocrystaline mercury sulfide thin films growth optical band gap optical property scanning electron microscopy sem successive ionic layer adsorption and reaction successive ionic layer adsorption and reaction trend transmission electron microscopy x ray diffraction dc two point probe method diameter respectively electrical properties electrical resistivity fine crystallites fine nanocrystalline film growth grown onto amorphous glass substrate immersion cycles optical band gap eg reaction silar trend room temperature c smooth surface morphology structural surface morphological optical successive ionic layer adsorption tem images confirmed films
Supplementary Fig. 7: Concept Extraction for W1993498556 (correct interpretation of ’and’; abbreviation not extracted).
S2 Concept graph properties
Supplementary Table 1: 25 most frequent concepts with their occurrences and degree in the concept graph.
@@
Supplementary Fig. 8: Distribution of node degrees. The node degree of 35 is the most common, occurring 2637 times. The concept with the highest degree is ’mechanical property’. The concept with the lowest degree is ’flexible pneumatic cylinder’. Both axes are in log scale.
Supplementary Fig. 9: Concept Centralization over different years. The graph illustrates the cumulative percentage of edges connected to nodes when sorted by their degree. A steep curve indicates that a small percentage of nodes (highly central nodes) are connected to a large percentage of edges. The years 1980, 1990, 2000, 2010, and $20$ are compared to show how concept centralization increases over time in our Concept Graph.
Supplementary Table 2: 5 exemplary concepts and their respective 5 nearest neighbors
@@
S3 Link prediction
Test set construction
To get a clear picture of our models’ performances, we do not perform oversampling for the positives during test set construction. This is necessary to analyze AUROC, precision and recall in a setting as close to real-world applications as possible. We approach test set construction by randomly sampling node pairs that are not connected until the end of 2019 and simply storing the label by looking three years ahead in the graph. The resulting 2,000,000 node pairs, therefore, mirror the real distribution of positives. Of these 307 (0.015%) positive examples, 87 links were formed in the year 2020, 123 in 2021, and 97 in 2022.
Evaluation against Science4Cast
We benchmark our Baseline model against the Science4Cast benchmark described in [14]. First, we train the model on data up until 2014 to predict links emerging in 2015, 2016 and 2017. Second, we evaluate on the node pairs contained in ‘SemanticGraph_delta_3_cutoff_0_minedge_1.pkl’. Surprisingly, our Baseline model is not only competitive but would have scored the second place in the Science4Cast competition with an AUROC of $0.9088$. Supplementary Table 3 presents an exhaustive overview of precision, recall, F1 score and the raw data at different thresholds.
Supplementary Table 3: Performance metrics of our Baseline model at different thresholds evaluated against the Science4Cast benchmark on ‘SemanticGraph_delta_3_cutoff_0_minedge_1.pkl’ [14].
Threshold AUC Precision Recall F1-score TN FP FN TP
0.50 0.9088 0.0818 0.5170 0.1413 9663029 287423 23934 25614
0.55 0.9088 0.0952 0.4705 0.1583 9728873 221579 26238 23310
0.60 0.9088 0.1101 0.4266 0.1751 9779674 170778 28411 21137
0.65 0.9088 0.1267 0.3837 0.1905 9819397 131055 30538 19010
0.70 0.9088 0.1458 0.3410 0.2042 9851432 99020 32653 16895
0.75 0.9088 0.1693 0.2950 0.2152 9878748 71704 34930 14618
0.80 0.9088 0.2027 0.2497 0.2238 9901799 48653 37175 12373
0.85 0.9088 0.2474 0.1975 0.2196 9920681 29771 39763 9785
0.90 0.9088 0.3225 0.1332 0.1885 9936587 13865 42948 6600
0.95 0.9088 0.4881 0.0568 0.1017 9947501 2951 46734 2814
Analysis of Ensemble Performance Across Feature Domains
We investigated the impact of model blending on AUC performance, i.e. whether performance gains arise from combining distinct feature modalities or solely from the averaging effect of ensembling. While blends of models with complementary information (embeddings mixed with topological models, Panels 1 and 2 of Supplementary Fig. 10) achieve substantial performance boosts, blends within the same feature domain (Panel 3) or model type (Panel 4) conversely result in only marginal improvements. This demonstrates that the observed gains are driven by the synergy of non-redundant features rather than simple ensemble averaging.
Supplementary Fig. 10: AUC performance for four hybrid models as a function of the blend weights for different combinations of models.
Calibration of the models’ predictions
Supplementary Table 4: Calibration of our models’ predictions measured with Brier-score.
Model Brier Score
Baseline 0.0521
GNN 0.0953
Concept Embeddings (MatSciBERT) 0.0535
Concept Embeddings (BERT) 0.0596
Combination of features 0.05019
Mixture of MLP and Embeddings 0.0438
Mixture of GNN and Embeddings 0.0438
Pure Text Baseline 0.1044
Precision and recall
Supplementary Fig. 11: Precision-recall (PR) curves for our link prediction models on the test set ($T_{test}=[2020,2022]$) with a zoom on the y-axis. Markers highlight the performances at a threshold of 0.5. The left panel shows the PR curve for all data points. The central and right panel display the respective PR curves for $d_{\text{prev}}=2$ and $d_{\text{prev}}=3$. Best result is in bold.
Supplementary Fig. 12: Precision@k for $k\in\{10,100,1000\}$ for our link prediction models for all data points, also broken down by year (2020, 2021, 2022).
Supplementary Fig. 13: Recall@k for $k\in\{10,100,1000\}$ for our link prediction models for all data points, also broken down by year (2020, 2021, 2022).
Supplementary Fig. 14: Precision@k for $k\in\{10,100,1000\}$ for our link prediction models for $d_{\text{prev}}=3$, also broken down by year (2020, 2021, 2022).
Supplementary Fig. 15: Recall@k for $k\in\{10,100,1000\}$ for our link prediction models for $d_{\text{prev}}=3$, also broken down by year (2020, 2021, 2022).
S4 Report generation
Supplementary Fig. 16: Schematic representation of the report structure. ’Own’ ($C_{\text{own}}$) and ’Other’ ($C_{\text{other}}$) concepts are respectively represented in yellow and blue.
@@
$S^{\text{filtered}}_{\text{own}\times\text{other}}:=\{$ $(c_{\text{own}},c_{{\text{other}}})\in Combinations_{\text{own},{\text{other}}}$
$|$ $\deg(c_{{\text{other}}})<=100\And 3<=\text{dist}(c_{\text{own}},c_{{\text{other}}}<=6)\}\quad.$
To avoid unspecific suggestions, i.e. broad terms like ‘materials science’, we exclude concepts based on their degree in the concept graph $G$. To prevent obvious, e.g. synonyms, and too far-fetched concepts, we integrate semantic knowledge by using the euclidean distance in the map of the materials science (Figure 1) that is a two dimensional projection of the obtained concept embeddings. Note that the values 3 and 6 were arbitrarily chosen after examining the map of materials science.
@@
$S_{\text{(many own)}\times\text{other}}$ $:=\{(c_{{\text{other}}},C_{\text{own},\text{matching}})\in C_{{\text{other}}}\times\{C_{\text{own}}\times...\times C_{\text{own}}\}|C_{\text{own},\text{matching}}$
$:=\{c\in C_{\text{own}}|\text{score}(c_{{\text{other}}},c)>p\}\}$
@@
We are interested in the capabilities of LLMs to assist humans by pre-selecting from a broader list of concept combinations. For this, we query two LLMs, GPT-4 and Claude 3 Opus (access May/June 2024), to select three items from a list of combinations and to write an additional paragraph on why this combination is promising. Three combinations were selected from $S_{\text{own}\times\text{own}}\cup S_{\text{own}\times\text{other}}$ together with three additional combinations from $S^{\text{filtered}}_{\text{own}\times\text{other}}$, resulting in a total of six suggested combinations per LLM per report, and the two prompts are given in s 17 and 18.
No paragraphs were created for combinations from section $S_{\text{(many own)}\times\text{other}}$ as we expected that having the LLM generate a full paragraph combining multiple concepts would be harder to digest and analyze within the limited interview time. Note that this report section does not contain new suggestions but rather selects and elaborates on combinations already present in the report. Apart from highlighting particular suggestions, the LLM approach helps to bridge the gap between two concepts by providing additional information on how they could be combined. This is especially helpful for $S_{\text{own}\times\text{other}}$ or $S^{\text{filtered}}_{\text{own}\times\text{other}}$ where the second concept might be unfamiliar to a researcher.
As our prompt to select 3 combinations of concepts from section 3 (Supplementary Fig. 18) explicitly asked the LLM to choose “more exotic” combinations, we analyzed if this resulted in a bias towards far-distance, cross-field pairs. For Sections 1 and 2, the mean node degree of the selected pairs was slightly lower (9,712) than for the non-selected pairs (11718). While all of the selected pairs were previously connected through only one additional node ($d_{\text{prev}}=2$), 2 of the non-selected pairs showed a larger previous node distance of $d_{\text{prev}}=3$. For section 3, mean node degrees of 1,582 and 4,694 were found for the selected and non-selected pairs, respectively. Here, 4 of the selected and 3 of the non-selected pairs showed a previous node distance of $d_{\text{prev}}=3$. While the lower node degrees as well as the slightly higher number of more distant concepts selected from section 3 could result from the wording in the prompt, the small sample size prevents any definitive conclusions.
Supplementary Table 5: LLM suggestion vs. "Is Interesting" (labeled as C) confusion matrix.
LLM suggestion
Y N $\sum$
"Is Interesting" Y 24 37 61
N 29 24 205
$\sum$ 53 213 266
You are an advanced AI agent who has expertise in the realm of the materials sciences. Your goal is to assist researchers looking for promising new concept combinations that will --- once investigated --- advance mankind.
@@
Supplementary Fig. 17: Prompt for the selection of 3 concept combinations from report sections 1 ($S_{\text{own}\times\text{own}}$) and 2 ($S_{\text{own}\times\text{other}}$).
@@
Supplementary Fig. 18: Prompt for the selection of 3 concept combinations from report section 3 ($S^{\text{filtered}}_{\text{own}\times\text{other}}$).
@@
Supplementary Fig. 19: Example of an LLM output containing three concept combinations and the corresponding abstracts
S5 Rated Concept Combinations
Selection of the interviewees
The selection of 10 interviewed researchers included only PIs, mostly professors at universities, from a diverse set of subdisciplines of materials science, some of which, but not all, had previous research experience with machine learning models.
Analysis of the rated concepts per researcher
Supplementary Fig. 20: Overview of the concept categories per interview. A1 and A2 are grouped together. Note that the results from researcher 3 were obtained by a commentary on the report, not by an interview.
Supplementary Table 6: Category counts per researcher with per-participant variance.
Researcher A1 A2 B C D Variance
1 11 9 19 9 0 36.64
2 3 7 5 4 0 5.36
3 9 0 1 12 1 24.24
4 18 2 7 3 0 41.20
5 8 7 15 9 0 22.96
6 4 3 11 11 1 17.60
7 4 0 9 8 2 11.84
8 3 3 6 6 0 5.04
9 5 4 23 7 5 51.36
10 6 1 3 8 0 9.04
Examples of rated concept combinations
Supplementary Table 7: Concept combinations ranked A1 and A2 during the interviews (25 examples per category)
@@
Supplementary Table 8: Concept combinations ranked A1 and A2 during the interviews (25 examples per category)
Category Concept A Concept B Section[a]
@@
ferroelectric domain interface bonding 3
@@
oxygen vacancy crack tip plasticity 4
stress induced phase transformation hexagonal boron nitride 3
D power consumption nanoparticle formation 3
quantum efficiency additive manufacturing 2
bauschinger effect ultimate tensile strength 2
moores law high entropy alloy 1